Figure 1: Behavior of the production and decay modes (see Eq.(14)) to the most relevant channels at present and future collider in three different SUSY breaking scenarios. When there is suppression or enhancement we indicate its maximum magnitude together with the parameters where this happens. If not stated explicitly, for the rest of the parameter space the mode behaves like the SM. The Higgs mass range where detection of a statistically signal is possible is also shown. For the LHC, the numbers follow the Ref.[hep-ph/0203123] and for LC and μC the Ref.[hep-ph/9703381].